OpenStudy (anonymous):

Given that f(x) = 3x + 1 and g(x) = the quantity of 4x plus 2 over 3, solve for g(f(0)). 2 6 8 9

OpenStudy (anonymous):

\[\frac{ 4x + 2 }{ 3 }\]

OpenStudy (solomonzelman):

4x+2 ---- 3 is the g(x) to find f(g(x)), plug in what the f(x) equals for x, INTO g(x). Can you do that ?

OpenStudy (solomonzelman):

I mean that would be g(f(x)), not f(g(x))

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ g(x)=\frac{4x+2}{3} }\) \(\LARGE\color{black}{ g(f(x))=\frac{4(3x+1)+2}{3}= \frac{12x+4+2}{3}= \frac{12x+6}{3} }\) \(\LARGE\color{black}{ = \frac{3(4x+2)}{3}=4x+2 }\)

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ g(f(0))=4(0)+2=?}\)

OpenStudy (anonymous):

sorry was afk. \[g(f(0)) = 4(0) + 2 = 2? \]

OpenStudy (anonymous):

Ok so A. these type of functions have always been difficult to me, thanks for the help